The rate constant of a first-order reaction is 3.9
Answer #1

Using the arrhenius equation

ln(k2/k1) = Ea/R * (1/T1 - 1/T2)

substituting the values in the equation we get

ln(7.30/3.90) = 131000/8.314 * (1/(350+273) - 1/T2)

1/T2 = 0.00156534

T2 = 638.83K = 365.83C

Hence the temperature in celsius scale will be equal to 365.83C